i need help with finding the length and the width with this question; The perimeter of a rectangle is 120 m. The Length is 9 m more than twice the width. find the dimensions.
Answer by checkley77(12844)

P=2L+2W
L=2W+9
120=2(2W+9)+2W
120=4W+18+2W
120=6W+18
6W=120-18
6W=102
W=102/6
W=17 ANS. FOR THE WIDTH.
L=2*17+9=34+9=43 ANS. FOR THE LENGTH.
PROOF:
120=2*43+2*17
120=86+34
120=120